# A Radio Map Approach for Reduced Pilot CSI Tracking in Massive MIMO Networks

### v3

@@ A Radio Map Approach for Reduced Pilot CSI Tracking in Massive MIMO Networks
Yuanshuai Zheng and Junting Chen Junting Chen is the corresponding author. The work was supported in part by the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone, by the National Science Foundation of China under Grant No. 62171398 and No. 62293482, by the Shenzhen Science and Technology Program under Grant No. JCYJ20220530143804010, No. KQTD20200909114730003, and No. KJZD20230923115104009, by Guangdong Research Projects No. 2019QN01X895, No. 2017ZT07X152, and No. 2019CX01X104, by the Shenzhen Outstanding Talents Training Fund 202002, by the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), by the National Key R&D Program of China with grant No. 2018YFB1800800, and by the Key Area R&D Program of Guangdong Province with grant No. 2018B030338001. School of Science and Engineering, Shenzhen Future Network of Intelligence Institute (FNii-Shenzhen), and
@@
Massive multiple-input multiple-output (MIMO) systems offer significant potential to enhance wireless communication performance, yet accurate and timely channel state information (CSI) acquisition remains a key challenge. Existing works on CSI estimation and radio map applications typically rely on stationary CSI statistics and accurate location labels. However, the CSI process can be discontinuous due to user mobility and environmental variations, and inaccurate location data can degrade the performance. By contrast, this paper studies radio-map-embedded CSI tracking and radio map construction without the assumptions of stationary CSI statistics and precise location labels. Using radio maps as the prior information, this paper develops a radio-map-embedded switching Kalman filter (SKF) framework that jointly tracks the location and the CSI with adaptive beamforming for sparse CSI observations under reduced pilots. For radio map construction without precise location labels, the location sequence and the channel covariance matrices are jointly estimated based on a Hidden Markov Model (HMM). An unbiased estimator on the channel covariance matrix is found. Numerical results on ray-traced MIMO channel datasets demonstrate that using $1$ pilot in every $10$ milliseconds, an average of over $97\%$ of capacity over that of perfect CSI can be achieved, while a conventional Kalman filter (KF) can only achieve $76\%$. Furthermore, the proposed radio-map-embedded CSI model can reduce the localization error from $30$ meters from the prior to $6$ meters for radio map construction.
@@ I Introduction
  • Can radio map help reduce pilots for CSI tracking? In particular, we focus on the case where there is no precise location information available as an index for the CSI in the radio map. Instead, location also needs to be recovered from the pilots as an intermediate step for the CSI tracking. Thus, we aim at a joint recovery of the location and CSI using radio maps as the prior information.
@@
  • Can we reconstruct and update a radio map from pilot sequences without precise location labels? To construct a radio map, we need to recover precise location information from pilot sequences with possibly some side information such as the locations of base stations (BSs). Thus, the question is, can a radio-map-embedded CSI model help recover precise location information despite possibly NLOS?
@@
This paper exploits the principle that while the CSI process can be discontinuous in a high-dimensional CSI space, the trajectory of the correspondingly mobile terminal has to be continuous in the three-dimensional physical world. As such, a sequence of very sparse CSI observations may suffice to discover the mobile trajectory, which leads to accurate tracking of the high-dimensional CSI with the help of radio maps. Following such a methodology, we develop a radio-map-embedded switching Kalman filter (SKF) framework that jointly tracks the location and the CSI with adaptive beamforming for sparse CSI observations under reduced pilots, where radio maps play an essential role in adjusting the operating regime of SKF. In addition, we employ a Bayesian approach for a joint estimation of the location and the channel covariance matrix from a massive amount of offline unlabeled pilot sequences for the construction of radio maps. The key theoretical contributions and numerical findings are summarized as follows:
@@
  • We propose a radio-map-embedded SKF framework for CSI tracking from a sequence of reduced pilot CSI measurements without location labels. In addition, instead of using a randomized compressive sensing approach, we develop an adaptive sensing matrix to minimize the differential entropy of the CSI.
@@
  • We develop a joint estimation on the location sequence and the channel covariance matrices from unlabeled pilot sequences using a Hidden Markov Model (HMM). An unbiased estimator on the channel covariance matrix is found with the estimation error analyzed.
@@
  • We numerically demonstrate that, in a single-user $64$-antenna Multiple-Input Single-Output (MISO) system, using $1$ pilot in every $10$ milliseconds, an average of over $97\%$ of capacity over that of perfect CSI can be achieved for a user moving at $36$ km/h at a $20$ dB signal-to-noise ratio (SNR) under LOS, while a conventional KF can only achieve $76\%$. In a pure NLOS regime, the proposed radio-map-assisted approach can double the capacity from existing schemes under low to moderate SNR. For radio map construction with unlabeled pilot sequences, we demonstrate that the proposed estimation using the radio-map-embedded CSI model reduces the localization error from $30$ meters from the prior to $6$ meters, which is sufficient for radio map construction in our framework.
@@
Notation: Vectors and matrices are denoted by bold lowercase $\mathbf{x}$ and bold uppercase $\mathbf{X}$, respectively. Scalar quantities are denoted by non-bold letters such as $x$. Collections are denoted by script $\mathcal{X}$. Probability density and mass functions are denoted by $p(\cdot)$ and $\mathbb{P}(\cdot)$, respectively. The expectation and variance operators are represented by $\mathbb{E}(\cdot)$ and $\mathbb{V}(\cdot)$, respectively. The $L_{2}$ norm is denoted by $\|\cdot\|_{2}$, and the Frobenius norm is denoted by $\|\cdot\|_{{\rm F}}$. The conjugate transpose of a complex vector $\mathbf{x}$ or a complex matrix $\mathbf{X}$ is denoted by ${\bf x}^{{\rm H}}$ or ${\bf X}^{{\rm H}}$, respectively. The trace of a matrix ${\bf X}$ is denoted by ${\rm tr}({\bf X})$. The identity matrix is denoted by ${\bf I}$. Furthermore, we specify several key variables that will be used throughout the paper: the user position, channel state, sensing matrix, and observation at time $t$ are denoted by $\mathbf{p}_{t}$, $\mathbf{h}_{t}$, $\mathbf{A}_{t}$, and $\mathbf{y}_{t}$, respectively. Their corresponding sequences up to time $t$ are denoted by $\mathcal{P}_{t}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{t})$, $\mathcal{H}_{t}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{t})$, $\mathcal{A}_{t}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{t})$, and $\mathcal{Y}_{t}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$t$}})$, with their realizations when $t=T$ denoted by $\mathcal{P}_{T}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{T})$, $\mathcal{H}_{T}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{T})$, $\mathcal{A}_{T}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T})$, and $\mathcal{Y}_{T}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$T$}})$, respectively.
@@ II-A Network Topology, Channel Model, and Radio Map Model
${\bf h}(\mathbf{p})=\sum_{l\in\mathcal{N}_{\text{s}}(\mathbf{p})}a_{l}(\mathbf{p})\bm{\alpha}(\theta_{l}(\mathbf{p}))+\sum_{m\in\mathcal{N}_{\text{m}}(\mathbf{p})}\beta_{m}\bm{\alpha}(\varphi_{m}),$ (1)
@@
${\bf h}(\mathbf{p})\triangleq{\bf h}^{\text{s}}(\mathbf{p})+{\bf h}^{\epsilon},$ (2)
@@
Consider a discrete location model, where the area of interest is partitioned into equally spaced grids. The user is located at one of the grid points $\mathbf{x}\in\mathcal{X}$, where $\mathcal{X}$ is a collection of all the grid locations. We define the radio map as a collection of the discretized location $\mathbf{x}\in\mathcal{X}$ and the covariance matrix ${\bf C}({\bf x})$ pairs, i.e., =$\mathcal{M}=\{(\mathbf{x},{\bf C}({\bf x})):\mathbf{x}\in\mathcal{X}\}$, where $\mathbf{C}(\mathbf{x})\triangleq\mathbb{E}\{\mathbf{h}\mathbf{h}^{\text{H}}\}$, in which the expectation is taken over the distribution of the positions $\mathbf{p}$ within the grid cell with a center location $\mathbf{x}\in\mathcal{X}$ and the small-scale fading of the channel. We assume that ${\bf h}^{\text{s}}(\mathbf{p})$ is spatially correlated for all the positions $\mathbf{p}$ in the same grid cell and $\mathbb{E}\{{\bf h}^{\epsilon}({\bf h}^{\epsilon})^{\text{H}}\}=\sigma_{\text{h}}^{2}\mathbf{I}$, where $\sigma_{\text{h}}^{2}$ is much smaller than the mean energy of ${\bf h}^{\text{s}}(\mathbf{p})$. Hence, the distribution of the channel $f_{\mathrm{h}}(\mathbf{h})$ given the location $\mathbf{x}$ is fully captured by the covariance matrix ${\bf C}({\bf x})$ and the parameter $\sigma_{\mathrm{h}}^{2}$. With such a discretized model, we now consider that $\mathbf{p}$ takes a discrete position from $\mathcal{X}$ in the rest of the paper.
@@ II-B System Dynamics
$\mathbb{P}(\mathbf{p}_{\text{$t$}}={\bf x}_{i}|\mathbf{p}_{t-1}={\bf x}_{j})\\ \propto\begin{cases}\exp(-\|{\bf x}_{i}-({\bf x}_{j}+\delta_{t}\bar{{\bf v}})\|_{2}^{2})&\text{if }\|{\bf x}_{i}-{\bf x}_{j}\|_{2}^{2}\leq\eta\delta_{t}\\ 0&\text{otherwise,}\end{cases}$ (3)
@@
A radio-map-embedded first-order AR model [8] for the dynamic of the channel $\mathbf{h}_{t}$ is considered as
${\bf h}_{t}=\gamma{\bf h}_{t-1}+\sqrt{1-\gamma^{2}}{\bf u}_{t},$ (4)
@@
The observation model (5) provides a unified representation for both uplink and downlink implementations. In the uplink case, the BS receives the signal ${\bf y}_{t}={\bf A}_{t}{\bf h}_{t}s+{\bf n}_{t}\in\mathbb{C}^{M}$ through the beamforming matrix ${\bf A}_{t}$, where the rows serve as combining vectors. The parameter $M$ may not be greater than the number of available radio-frequency (RF) chains at the BS. We can simply take $s=1$ without loss of generality (w.l.o.g.) to yield the observation model (5). In the downlink case, the BS transmits a beamformed pilot ${\bf a}_{t}^{*}s$ (with $s$ again assumed to be $1$) where ${\bf a}_{t}^{*}$ is the conjugate of $\mathbf{a}_{t}$, and the user receives the signal as $y_{t}={\bf h}_{t}^{{\rm T}}{\bf a}_{t}^{*}s+n_{t}$, where $y_{t}$ and $n_{t}$ denote the scalar observation and noise, respectively. Then, the user may feed back the estimated channel or simply the received scalar $y_{t}$ to the BS. The matrix ${\bf A}_{t}$ in this case comprises $M$ beamforming vectors, with each row of ${\bf A}_{t}$ given by the conjugate transpose of a beamforming vector ${\bf a}_{t}$. Thus, the model in (5) also captures the downlink scenario in a unified form.

### v5

@@ A Radio Map Approach for Reduced Pilot CSI Tracking in Massive MIMO Networks
Yuanshuai Zheng and Junting Chen This is a preprint version. The paper has been published at IEEE Transactions on Signal Processing, vol. 73, pp. 2833-2847, 2025, doi: 10.1109/TSP.2025.3584229. Junting Chen is the corresponding author. The work was supported in part by the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone, by the National Science Foundation of China under Grant No. 62171398 and No. 62293482, by the Guangdong Basic and Applied Basic Research Foundation 2024A1515011206, by the Shenzhen Science and Technology Program under Grant No. JCYJ20220530143804010, No. KQTD20200909114730003, and No. KJZD20230923115104009, by Guangdong Research Projects No. 2019QN01X895, No. 2017ZT07X152, and No. 2019CX01X104, by the Shenzhen Outstanding Talents Training Fund 202002, by the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), by the National Key R&D Program of China with grant No. 2018YFB1800800, and by the Key Area R&D Program of Guangdong Province with grant No. 2018B030338001. School of Science and Engineering, Shenzhen Future Network of Intelligence Institute (FNii-Shenzhen), and
@@
Massive multiple-input multiple-output (MIMO) systems offer significant potential to enhance wireless communication performance, yet accurate and timely channel state information (CSI) acquisition remains a key challenge. Existing works on CSI estimation and radio map applications typically rely on stationary CSI statistics and accurate location labels. However, the CSI process can be discontinuous due to user mobility and environmental variations, and inaccurate location data can degrade the performance. By contrast, this paper studies radio-map-embedded CSI tracking and radio map construction without the assumptions of stationary CSI statistics and precise location labels. Using radio maps as the prior information, this paper develops a radio-map-embedded switching Kalman filter (SKF) framework that jointly tracks the location and the CSI with adaptive beamforming for sparse CSI observations under reduced pilots. For radio map construction without precise location labels, the location sequence and the channel covariance matrices are jointly estimated based on a Hidden Markov Model (HMM). An unbiased estimator on the channel covariance matrix is found. Numerical results on ray-traced MIMO channel datasets demonstrate that using $1$ pilot observation in every $10$ milliseconds, an average of over $97\%$ of capacity over that of perfect CSI can be achieved, while a conventional Kalman filter (KF) can only achieve $76\%$. Furthermore, the proposed radio-map-embedded CSI model can reduce the localization error from $30$ meters from the prior to $6$ meters for radio map construction.
@@ I Introduction
  • Can radio map help reduce pilot observations for CSI tracking? In particular, we focus on the case where there is no precise location information available as an index for the CSI in the radio map. Instead, location also needs to be recovered from the pilot observations as an intermediate step for the CSI tracking. Thus, we aim at a joint recovery of the location and CSI using radio maps as the prior information.
@@
  • Can we reconstruct and update a radio map from pilot observation sequences without precise location labels? To construct a radio map, we need to recover precise location information from pilot observation sequences with possibly some side information such as the locations of base stations (BSs). Thus, the question is, can a radio-map-embedded CSI model help recover precise location information despite possibly NLOS?
@@
In contrast to the preliminary study [14] which maps a location to a specific channel realization, this work adopts a different radio map methodology, where it defines a mapping from a location to a channel distribution. As such, the new radio map definition is robust to local scattering, body shadowing, and antenna pattern at the user side. Consequently, it enables not only robust radio-map-embedded CSI tracking, but also a blind construction of radio maps without location labels. Towards this goal, this paper exploits the principle that while the CSI process can be discontinuous in a high-dimensional CSI space, the trajectory of the correspondingly mobile terminal has to be continuous in the three-dimensional physical world. As such, a sequence of very sparse CSI observations may suffice to discover the mobile trajectory, which leads to accurate tracking of the high-dimensional CSI with the help of radio maps. Following such a methodology, we develop a radio-map-embedded switching Kalman filter (SKF) framework that jointly tracks the location and the CSI with adaptive beamforming for sparse CSI observations under reduced pilot observations, where radio maps play an essential role in adjusting the operating regime of SKF. In addition, we employ a Bayesian approach for a joint estimation of the location and the channel covariance matrix from a massive amount of offline unlabeled pilot observation sequences for the construction of radio maps. The key theoretical contributions and numerical findings are summarized as follows:
@@
  • We propose a radio-map-embedded SKF framework for CSI tracking from a sequence of reduced pilot observations without location labels. In addition, instead of using a randomized compressive sensing approach, we develop an adaptive sensing matrix to minimize the differential entropy of the CSI.
@@
  • We develop a joint estimation on the location sequence and the channel covariance matrices from unlabeled pilot observation sequences using a Hidden Markov Model (HMM). An unbiased estimator on the channel covariance matrix is found with the estimation error analyzed.
@@
  • We numerically demonstrate that, in a single-user $64$-antenna Multiple-Input Single-Output (MISO) system, using $1$ pilot observation in every $10$ milliseconds, an average of over $97\%$ of capacity over that of perfect CSI can be achieved for a user moving at $36$ km/h at a $20$ dB signal-to-noise ratio (SNR) under LOS, while a conventional KF can only achieve $76\%$. In a pure NLOS regime, the proposed radio-map-assisted approach can double the capacity from existing schemes under low to moderate SNR. For radio map construction with unlabeled pilot observation sequences, we demonstrate that the proposed estimation using the radio-map-embedded CSI model reduces the localization error from $30$ meters from the prior to $6$ meters, which is sufficient for radio map construction in our framework.
@@
Notation: Vectors and matrices are denoted by bold lowercase $\mathbf{x}$ and bold uppercase $\mathbf{X}$, respectively. Scalar quantities are denoted by non-bold letters such as $x$. Collections are denoted by script $\mathcal{X}$. Probability density and mass functions are denoted by $p(\cdot)$ and $\mathbb{P}(\cdot)$, respectively. The expectation and variance operators are represented by $\mathbb{E}(\cdot)$ and $\mathbb{V}(\cdot)$, respectively. The $L_{2}$ norm is denoted by $\|\cdot\|_{2}$, and the Frobenius norm is denoted by $\|\cdot\|_{{\rm F}}$. The conjugate transpose of a complex vector $\mathbf{x}$ or a complex matrix $\mathbf{X}$ is denoted by ${\bf x}^{{\rm H}}$ or ${\bf X}^{{\rm H}}$, respectively. The trace of a matrix $\bf X$} is denoted by ${\rm tr}({\bf X})$. The pseudo-inverse of a matrix $\mathbf{X}$ is denoted by $\mathbf{X}^{\dagger}$. The identity matrix is denoted by $\bf I$}. Furthermore, we specify several key variables that will be used throughout the paper: the user position, channel state, sensing matrix, and observation at time $t$ are denoted by $\mathbf{p}_{t}$, $\mathbf{h}_{t}$, $\mathbf{A}_{t}$, and $\mathbf{y}_{t}$, respectively. Their corresponding sequences up to time $t$ are denoted by $\mathcal{P}_{t}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{t})$, $\mathcal{H}_{t}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{t})$, $\mathcal{A}_{t}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{t})$, and $\mathcal{Y}_{t}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$t$}})$, with their realizations when $t=T$ denoted by $\mathcal{P}_{T}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{T})$, $\mathcal{H}_{T}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{T})$, $\mathcal{A}_{T}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T})$, and $\mathcal{Y}_{T}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$T$}})$, respectively.
@@ II-A Network Topology, Channel Model, and Radio Map Model
${\bf h}(\mathbf{p})=\sum_{l\in\mathcal{N}_{\text{s}}(\mathbf{p})}a_{l}(\mathbf{p})\bm{\alpha}(\theta_{l}(\mathbf{p}))+\sum_{m\in\mathcal{N}_{\text{m}}(\mathbf{p})}\beta_{m}\bm{\alpha}(\varphi_{m})$ (1)
@@
${\bf h}(\mathbf{p})\triangleq{\bf h}^{\text{s}}(\mathbf{p})+{\bf h}^{\epsilon}$ (2)
@@
Consider a discrete location model, where the area of interest is partitioned into equally spaced grids. The user is located at one of the grid points $\mathbf{x}\in\mathcal{X}$, where $\mathcal{X}$ is a collection of all the grid locations. We define the radio map as a collection of the discretized location $\mathbf{x}\in\mathcal{X}$ and the covariance matrix ${\bf C}({\bf x})$ pairs, i.e., $\mathcal{M}=\{(\mathbf{x},{\bf C}({\bf x})):\mathbf{x}\in\mathcal{X}\}$, where $\mathbf{C}(\mathbf{x})\triangleq\mathbb{E}\{\mathbf{h}\mathbf{h}^{\text{H}}\}$, in which the expectation is taken over the distribution of the positions $\mathbf{p}$ within the grid cell with a center location $\mathbf{x}\in\mathcal{X}$ and the small-scale fading of the channel. We assume that ${\bf h}^{\text{s}}(\mathbf{p})$ is spatially correlated for all the positions $\mathbf{p}$ in the same grid cell and $\mathbb{E}\{{\bf h}^{\epsilon}({\bf h}^{\epsilon})^{\text{H}}\}=\sigma_{\text{h}}^{2}\mathbf{I}$, where $\sigma_{\text{h}}^{2}$ is much smaller than the mean energy of ${\bf h}^{\text{s}}(\mathbf{p})$. Hence, the distribution of the channel $f_{\mathrm{h}}(\mathbf{h})$ given the location $\mathbf{x}$ is fully captured by the covariance matrix ${\bf C}({\bf x})$ and the parameter $\sigma_{\mathrm{h}}^{2}$. With such a discretized model, we now consider that $\mathbf{p}$ takes a discrete position from $\mathcal{X}$ in the rest of the paper.
@@ II-B System Dynamics
$\mathbb{P}(\mathbf{p}_{\text{$t$}}={\bf x}_{i}|\mathbf{p}_{t-1}={\bf x}_{j})\\ \propto\begin{cases}\exp(-\|{\bf x}_{i}-({\bf x}_{j}+\delta_{t}\bar{{\bf v}})\|_{2}^{2})&\text{if }\|{\bf x}_{i}-{\bf x}_{j}\|_{2}^{2}\leq\eta\delta_{t}\\ 0&\text{otherwise}\end{cases}$ (3)
@@
A radio-map-embedded first-order AR model [8, 25] for the dynamic of the channel $\mathbf{h}_{t}$ is considered as
${\bf h}_{t}=\gamma{\bf h}_{t-1}+\sqrt{1-\gamma^{2}}{\bf u}_{t}$ (4)
@@
The observation model (5) provides a unified representation for both uplink and downlink implementations. In the uplink case, the BS receives the signal ${\bf y}_{t}={\bf A}_{t}{\bf h}_{t}s+{\bf n}_{t}\in\mathbb{C}^{M}$ through the beamforming matrix ${\bf A}_{t}$, where the rows of ${\bf A}_{t}$ serve as combining vectors and $s$ denotes the reference symbol. The parameter $M$ may correspond to the number of available radio-frequency (RF) chains at the BS. We can simply take $s=1$ without loss of generality (w.l.o.g.) to yield the observation model (5). Note that $\mathbf{n}_{t}$ is the equivalent noise vector after combining the noise from the $N_{\mathrm{t}}$ antennas using the semi-unitary matrix ${\bf A}_{t}\in\mathbb{C}^{M\times N_{\mathrm{t}}}$. If $\tilde{\mathbf{n}}_{t}$ denotes the $N_{\mathrm{t}}\times 1$ noise vector at the antennas, then $\mathbf{n}_{t}=\mathbf{A}_{t}\tilde{\mathbf{n}}_{t}$. Since each row of $\mathbf{A}_{t}$ is orthonormal, both the elements of $\mathbf{n}_{t}$ and those of $\tilde{\mathbf{n}}_{t}$ are independent and identically distributed (i.i.d.) with zero mean and the same variance. In the downlink case, consider using $M$ orthogonal time-frequency resources for channel probing at time slot $t$. For the $m$th time-frequency resource, a reference symbol $s$ is beamformed using a vector $\mathbf{a}_{t,m}$, where $\mathbf{a}_{t,m}^{\text{T}}$ corresponds to the $m$th row of $\mathbf{A}_{t}$. As a result, the user obtains a noisy observation $y_{t,m}=\mathbf{h}_{t}^{\text{T}}\mathbf{a}_{t,m}s+n_{t,m}$, and by stacking the $M$ observations $y_{t,m}$ into a vector $\mathbf{y}_{t}$, the received signal model is identical to (5). In this case, the user may need to feed back the received pilot observations $\mathbf{y}_{t}$ to the BS or it is the user that performs the channel tracking.
Note that both downlink and uplink transmissions involve feedback signaling, such as scheduling grants and CSI feedback as specified in the 5G standard. However, measuring the feedback overhead, which is strongly implementation dependent, falls outside the scope of this paper. Therefore, this paper intentionally considers only the $M$ pilot observations as the sole contributors to the received vector $\mathbf{y}_{t}$ for CSI tracking.